Atoms of CO 2 gas decay from upper to lower level with a spread in frequency is about10kHz. Calculate the line shape if the broadening mechanism is found Doppler broadening?
1
Expert's answer
2020-05-29T10:08:47-0400

The frequencies will be distributed according to Gaussian line shape:


"L=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{(f-f_0)^2}{2\sigma^2}},"

in this equation


"\sigma=\frac{f_0}{c}\sqrt{\frac{2kT\text{ ln}2}{m}}."

More information is needed to answer this question.